# Quantifying Decoherence

### v2

@@ Abstract
Quantum decoherence refers to the phenomenon when the interaction of a quantum system with its environment results in the degradation of quantum coherence. Decoherence is considered to be the most popular mechanism responsible for the emergence of classicality from quantum mechanics. The issue of formulating a measure of decoherence is addressed here. The approach taken here is that decoherence results from the entanglement of a quantum system with certain environment degrees of freedom, and quantifying this entanglement should yield the most natural measure of decoherence. A simple measure of decoherence is presented based on this notion, and it is examined for various example systems. The measure proves to be effective and is relatively straightforward to compute. In addition, a method has been proposed to measure decoherence in a Mach-Zehnder interferometer.
@@ I Introduction
Early works on decoherence mostly focused on the time scales over which decoherence occurs [9], which is interesting for addressing the question if the emergence of classicality is instantaneous or occurs over a time scale. With it also arose a need to quantify how classical a state becomes as it interacts with the environment. Lee [17] introduced a measure of “non-classicality” of a quantum state $\rho$, the quantity $\tau_{\rho}$, referred to as the “non-classical depth.” This measure was later used by several authors. The subject of quantifying decoherence has been previously reviewed [18, 19]. In more recent studies Zhang and Luo introduced an interesting decoherence measure in terms of the averaged Wigner-Yanase skew information [20], Fu and Luo approached quantifying decoherence via increase in classicality [21], Gundhi and Ulbricht introduced a quantifier named “decoherence kernel” [22]. It has been experimentally demonstrated that partially coherent electron wave packets can be constructed, which highlights the necessity to develop methods for quantifying decoherence [23].
@@ III.1 Entanglement between the system and environment
From the above discussion it is clear that decoherence is the result of the system getting entangled with the environment. Our approach is that the degree of decoherence can be quantified by quantifying the entanglement between the system and environment. In this combined system-environment aggregate, involving a lot of degrees of freedom, we consider a bipartition of the system-environment aggregate such that our system of interest forms one partition, and the rest of the environment forms the other. One can now ask, how much is the entanglement between the system and the rest of the environment. The generalized concurrence for this bipartition can be written as [25]
@@
where $\rho$ is the reduced density operator of the system, after tracing over the environment states. In order to get the degree of decoherence of the system after a time $t$, in the scenario described above, we insert (5) in (8) to get
@@
If the system states do not get correlated to the environment at all, it implies that $\mathcal{U}_{t}^{(k){\dagger}}\mathcal{U}_{t}^{(j)}=1$ for all $j,k$. In this situation $\mathcal{D}_{e}=\tfrac{n}{n-1}\big[1-\sum_{i=1}^{n}|c_{i}|^{4}-\sum_{j\neq k}|c_{j}|^{2}|c_{k}|^{2}\big]=0$. If the environment states correlated to the states $|p_{i}\rangle$ are all orthogonal to each other, and $|c_{i}|^{2}=\tfrac{1}{n}$ for all $i$, $\mathcal{D}_{e}=\tfrac{n}{n-1}[1-\sum_{i=1}^{n}|c_{i}|^{4}]=1$. This is the case of maximal decoherence of the system. Here the reduced density operator of the system is maximally mixed.
@@ IV.2 Decoherence of a free particle
where $x,P_{x}$ represent the position and momentum operator of the particle, $r_{k},p_{k}$ are position and momentum of the k’th oscillator. This model was originally proposed for studying quantum dissipation or damping [37]. However, here we use it to study decoherence. The problem of a particle coupled to a a bath of independent harmonic oscillators was studied in detail by Caldeira and Leggett [38]. One can write a density operator for the particle and the heat-bath combined, and evolves it using the Hamiltonian (26). The dynamics of the particles can then be studied by averaging over the environment degrees of freedom, and looking at the reduced density operator of the particle. Under certain conditions the reduced density matrix of the particle follows the equation [38]
@@
where $\sigma=\frac{\pi}{\lambda_{T}^{2}}(1-e^{-4\gamma t})$. Before proceeding further we would like to emphasize the time scales one should look at while studying decoherence. In the problem studied by Savage and Walls [37], the parameter $\gamma$ has a simple meaning - it is the thermal relaxation rate of the particle due to the damping effect of the environment. However, while studying decoherence one is studying such minimal effects of the environment that no change in the energy of the particle results from it, and particle travels virtually undamped [36]. The effect of the environment is only a dephasing of its time evolution. Decoherence effects should come into play at times much shorter than $1/\gamma$ [36]. For $\gamma t\ll 1$ one can make the following approximation: $\sigma\approx\frac{4\pi}{\lambda_{T}^{2}}\gamma t$. One can now use (25) to evaluate the decoherence measure, and it comes out to be
@@
Notice that in the limit $\sqrt{\sigma}L\gg 1$, i.e., for $t\gg\frac{\lambda_{T}^{2}}{4\pi\gamma L^{2}}$ the decoherence measure $\mathcal{D}_{e}$ tends to its maximum value 1. The time scale that emerges here agrees with the decoherence time scale $\tau_{D}=\frac{\lambda_{T}^{2}}{4\pi\gamma L^{2}}$ estimated earlier [9]. For intermediate times one can estimate the degree of decoherence, the particle has undergone, from (29). The decoherence measure is plotted as a function of time (scaled by $\tau_{D}$) in Fig. 1. It is interesting to see that the growth of the degree of decoherence with time is distinctly slower than the naive expectation of $1-e^{-t/\tau_{D}}$.
@@
In the limit $t\to 0$, the decoherence measure $\mathcal{D}_{e}=1$. For $t\gg(\frac{m\sigma^{2}}{16k_{B}T\gamma})^{1/3}$ the decoherence measure approaches its maximal value 1. Interestingly, here the decoherence is not faster if the position spread of the particle is large, i.e., $\sigma$ is large. Rather, it is faster if the momentum spread of the particle is large, i.e., $\hbar/\sigma$ is large. This is probably an artifact of the coupling of the particle’s position to the momenta of the environment, in (26).
@@
A more realistic coupling to the environment is given by the Hamiltonian [38]
@@
The time evolution of the reduced density matrix is governed by the following master equation [38]
@@ IV.3 Decoherence in a Stern-Gerlach experiment
Next we look at an experiment which is a prototype for studying quantum measurement, namely, the Stern-Gerlach experiment. In this experiment we consider a spin-1/2 particle of mass $m$, traveling along y-axis, and passing through an inhomogeneous magnetic field pointing along z-axis. The field is inhomogeneous in the x-direction. Without going into calculational details, we use this setup to illustrate the usefulness the present decoherence measure in dealing with situations where there both discrete and continuous degrees of freedom. Two different kinds of decoherence effects are expected to happen in this experiment, and at two different time scales. It is interesting to see how the decoherence measure captures the two effects. If we just look at the evolution of the particle in the x-direction, the Hamiltonian can be written as [30]
@@ V Measuring decoherence
where $\rho_{11}$ ($\rho_{22}$) represents the probability of the quanton passing through the upper (lower) path. The decoherence measure can then be calculated using (8) as
@@
The problem at hand now is find a way to make measurements in the experiment to evaluate the above decoherence measure. If one blocks the lower path before it enters BS2, there will be equal intensity at D1 and D2 which will be proportional to $\rho_{11}$. On the other hand, if one blocks the upper path before it enters BS2, there will again be equal intensity at D1 and D2 which will now be proportional to $\rho_{22}$. It can be shown that the average intensity at the two detectors, in the two scenarios, is given by
@@
where we have assumed $\phi=0$. If one is unsure about $\phi$ in the setup, one can tune the phase shifter such that maximum value of $I_{diff}^{2}$ is obtained. In this analysis it is assumed that all intensities are normalized with the total intensity. Given that the Mach-Zehnder interferometer with cold neutrons is widely used and there have been significant developments in the field of neutron interferometry [39], the experiment suggested here should be feasible with neutrons, allowing for the measurement of decoherence. The investigation of decoherence through neutron interferometry has been previously suggested [40]. While decoherence in a conventional photonic Mach-Zehnder interferometer in a lab would be negligible, one may use the proposed method to measure decoherence in optical interferometry involving large distances, e.g., as in the setup involving gravitational wave detection. Interferometry with atoms and molecules is a topic of active research, and decoherence is an important issue there [41, 42, 43, 44]. One can think of extending the method proposed here to such matter-wave interferometry experiments.
@@ VI Discussion and conclusions
To summarize, we formulated a measure of decoherence based on the quantification of entanglement of a system with the environment. The measure is given by $\mathcal{D}_{e}=\tfrac{n}{n-1}[1-\text{tr}(\rho^{2})]$ for systems with finite-dimensional Hilbert space, and $\mathcal{D}_{e}=[1-\text{tr}(\rho^{2})]$ for continuous variable systems, where $\rho$ represents the reduced density operator of the system after tracing over the environment degrees of freedom. This decoherence measure has minimum value 0 and maximum value 1. Figuring out the pointer states that emerge out of decoherence of a system is often a difficult issue. In one basis, the off-diagonal elements might diminish more rapidly over time, whereas in another basis, they may also diminish, albeit at a slower rate [36]. The present formulation does not depend on the pointer states. One just needs to evaluate a trace which is basis independent.

### v3

@@ Abstract
Quantum decoherence refers to the phenomenon when the interaction of a quantum system with its environment results in the degradation of quantum coherence. Decoherence is considered to be the most popular mechanism responsible for the emergence of classicality from quantum mechanics. The issue of formulating a measure of decoherence is addressed here. The approach taken here is that decoherence results from the entanglement of a quantum system with certain environment degrees of freedom, and quantifying this entanglement should yield the most natural measure of decoherence. A simple measure of decoherence is presented based on this notion, and it is examined for various example systems. The measure proves to be effective and is relatively straightforward to compute. In addition, a method has been proposed to measure decoherence in a Mach-Zehnder interferometer which may be useful in neutron interferometry.
@@ I Introduction
Early works on decoherence mostly focused on the time scales over which decoherence occurs [9], which is interesting for addressing the question if the emergence of classicality is instantaneous or occurs over a time scale. With it also arose a need to quantify how classical a state becomes as it interacts with the environment. Lee [17] introduced a measure of “non-classicality” of a quantum state $\rho$, the quantity $\tau_{\rho}$, referred to as the “non-classical depth.” This measure was later used by several authors. The subject of quantifying decoherence has been previously reviewed [18, 19]. In more recent studies Zhang and Luo introduced an interesting decoherence measure in terms of the averaged Wigner-Yanase skew information [20], Fu and Luo approached quantifying decoherence via increase in classicality [21], Gundhi and Ulbricht introduced a quantifier named “decoherence kernel” [22]. It has been experimentally demonstrated that partially coherent electron wave packets can be constructed, which highlights the necessity to develop methods for quantifying decoherence [23]. In an interesting study, decoherence induced during the preparation of highly non-classical states for quantum metrology, arising from entanglement with a quantized auxiliary system was studied using quantum Fisher information (QFI) [24]. This approach applies to situations where the auxillary system cannot be probed, which is also the case with the environment in standard decoherence. However it seems more suited to specific scenarios like state preparation, and may not apply to, say, environment induced decoherence of a particle passing through a double slit. Nevertheless, QFI has become a popular tool to probe decoherence in many subsequent studies.
About a decade back, a specific measure of coherence was established, primarily driven by its necessity in the domain of quantum information [25]. This measure has demonstrated significant utility, even in quantifying the wave characteristics of quantum particles [26]. One might imagine that, given the strong association between decoherence and the loss of coherence, one could express decoherence in terms of the coherence measure. Nevertheless, the challenge in this approach lies in the fact that the measure of coherence is dependent on the basis, while it is desirable to have a measure of decoherence that is independent of the basis, for reasons that will become clear in the subsequent discussion.
As an open quantum system evolves, it typically establishes correlations with its environment and irreversibly diminishes its information content. Nevertheless, this holds true only when the system is weakly coupled to the environment. If the coupling of the system to the environment is strong, information may periodically revert to its original state from the environment, resulting in a back-flow of information. This phenomenon is referred to as non-Markovian behavior. Of late quantum non-Markovianity has become an active subject of research [27, 28, 29], and measures of the degree of non-Markovianity have been proposed [30, 31]. However, the decoherence induced emergence of classicality arises from an extremely weak interaction with the environment, so minimal that the system effectively evolves without dissipation. Consequently, non-Markovianity may not be applicable in these contexts. Therefore, it is necessary to establish a measure that is relevant to scenarios involving weak interactions with the environment.
@@ III.1 Entanglement between the system and environment
From the above discussion it is clear that decoherence is the result of the system getting entangled with the environment. Our approach is that the degree of decoherence can be quantified by quantifying the entanglement between the system and environment. In this combined system-environment aggregate, involving a lot of degrees of freedom, we consider a bipartition of the system-environment aggregate such that our system of interest forms one partition, and the rest of the environment forms the other. One can now ask, how much is the entanglement between the system and the rest of the environment. To this end we use the methodology developed by Bhaskara and Panigrahi [33] for quantifying multipartite entanglement. Let us assume that our system of interest has interacted with certain environment degrees of freedom, such that the combined state of the system and the environment is given by (4). We assume the the degrees of freedom described by the states $|p_{i}\rangle$ form one partition, and those described by the states $\mathcal{U}_{t}^{(i)}|\mathcal{E}^{0}\rangle$ form the other much larger partition. We wish to quantify the entanglement between these two partitions. In the methodology of Bhaskara and Panigrahi [33], one first traces over the degrees of freedom of one partition to obtain a reduced density operator. In our case, we trace over the environment states and get the reduced density operator $\rho$ for our system of interest. The generalized concurrence for this bipartition can then be written as [33]
@@
where $\rho$ is the reduced density operator of the system, after tracing over the environment states. We first test out this measure for the scenario described above. In order to get the degree of decoherence of the system after a time $t$, we plug in the reduced density operator from (5) in the expression for the decoherence measure (8). That procedure results in following form of the decoherence measure:
@@
If the system states do not get correlated to the environment at all, it implies that the environment states correlated to the system states, $\mathcal{U}_{t}^{(i)}|\mathcal{E}^{0}\rangle$, are all identical, and $\mathcal{U}_{t}^{(k){\dagger}}\mathcal{U}_{t}^{(j)}=1$ for all $j,k$. In this situation $\mathcal{D}_{e}=\tfrac{n}{n-1}\big[1-\sum_{i=1}^{n}|c_{i}|^{4}-\sum_{j\neq k}|c_{j}|^{2}|c_{k}|^{2}\big]=0$. If the environment states correlated to the states $|p_{i}\rangle$ are all orthogonal to each other, and $|c_{i}|^{2}=\tfrac{1}{n}$ for all $i$, $\mathcal{D}_{e}=\tfrac{n}{n-1}[1-\sum_{i=1}^{n}|c_{i}|^{4}]=1$. This is the case of maximal decoherence of the system. Here the reduced density operator of the system is maximally mixed.
@@ IV.2 Decoherence of a free particle
where $x,P_{x}$ represent the position and momentum operator of the particle, $r_{k},p_{k}$ are position and momentum of the k’th oscillator. This model was originally proposed for studying quantum dissipation or damping [46]. However, here we use it to study decoherence. The problem of a particle coupled to a bath of independent harmonic oscillators was studied in detail by Caldeira and Leggett [47, 48]. One can write a density operator for the particle and the heat-bath combined, and evolve it using the Hamiltonian (26). The dynamics of the particles can then be studied by averaging over the environment degrees of freedom, and looking at the reduced density operator of the particle. Under certain conditions the reduced density matrix of the particle follows the equation [47, 48]
@@
where $\sigma=\frac{\pi}{\lambda_{T}^{2}}(1-e^{-4\gamma t})$. Before proceeding further we would like to emphasize the time scales one should look at while studying decoherence. In the problem studied by Savage and Walls [46], the parameter $\gamma$ has a simple meaning - it is the thermal relaxation rate of the particle due to the damping effect of the environment. However, while studying decoherence one is studying such minimal effects of the environment that no change in the energy of the particle results from it, and particle travels virtually undamped [44, 45]. The effect of the environment is only a dephasing of its time evolution. Decoherence effects should come into play at times much shorter than $1/\gamma$ [44]. For $\gamma t\ll 1$ one can make the following approximation: $\sigma\approx\frac{4\pi}{\lambda_{T}^{2}}\gamma t$. One can now use (25) to evaluate the decoherence measure, and it comes out to be
@@
Notice that in the limit $\sqrt{\sigma}L\gg 1$, i.e., for $t\gg\frac{\lambda_{T}^{2}}{4\pi\gamma L^{2}}$ the decoherence measure $\mathcal{D}_{e}$ tends to its maximum value 1. Here $1/\sqrt{\sigma}$ is related to the width of the wavepacket at time $t$. So the limit implies the situation where decoherence is strong, and the wavepacket doesn’t spread much even after a time $t$. The time scale that emerges here agrees with the decoherence time scale $\tau_{D}=\frac{\lambda_{T}^{2}}{4\pi\gamma L^{2}}$ estimated earlier [9]. For intermediate times one can estimate the degree of decoherence, the particle has undergone, from (29). The decoherence measure is plotted as a function of time (scaled by $\tau_{D}$) in Fig. 1. It is interesting to see that the growth of the degree of decoherence with time is distinctly slower than the naive expectation of $1-e^{-t/\tau_{D}}$.
@@
In the limit $t\to 0$, the decoherence measure $\mathcal{D}_{e}=0$, implying that the environment influence is negligible during such short times. For longer times the environment has enough time to cause decoherence, so for $t\gg(\frac{m\sigma^{2}}{16k_{B}T\gamma})^{1/3}$ the decoherence measure approaches its maximal value 1. Interestingly, here the decoherence is not faster if the position spread of the particle is large, i.e., $\sigma$ is large. Rather, it is faster if the momentum spread of the particle is large, i.e., $\hbar/\sigma$ is large. This is probably an artifact of the coupling of the particle’s position to the momenta of the environment, in (26).
@@
A more realistic coupling to the environment is given by the Hamiltonian [47, 48]
@@
The time evolution of the reduced density matrix is governed by the following master equation [47, 48]
@@ IV.3 Decoherence in a Stern-Gerlach experiment
Next we look at an experiment which is a prototype for studying quantum measurement, namely, the Stern-Gerlach experiment. In this experiment we consider a spin-1/2 particle of mass $m$, traveling along y-axis, and passing through an inhomogeneous magnetic field pointing along z-axis. The field is inhomogeneous in the x-direction. Without going into calculational details, we use this setup to illustrate the usefulness of the present decoherence measure in dealing with situations where there are both discrete and continuous degrees of freedom. Two different kinds of decoherence effects are expected to happen in this experiment, and at two different time scales. It is interesting to see how the decoherence measure captures the two effects. If we just look at the evolution of the particle in the x-direction, the Hamiltonian can be written as [38]
@@ V Measuring decoherence
where $\rho_{11}$ ($\rho_{22}$) represents the probability of the quanton passing through the upper (lower) path. We would like to emphasize here that although the analysis started with a pure quanton state (44), purity of the initial state is not at all a requirement. The quanton may already be in a mixed state before entering the setup, but the expression (47) will still hold. The phase difference $\phi$ does not represent the phase coherence in the initial state. Rather the phase difference may be introduced in the two paths by the phase shifter shown in Fig. 3. The decoherence measure can then be calculated using (8) as
@@
The problem at hand now is to find a way to make measurements in the experiment to evaluate the above decoherence measure. If one blocks the lower path before it enters BS2, there will be equal intensity at D1 and D2 which will be proportional to $\rho_{11}$. On the other hand, if one blocks the upper path before it enters BS2, there will again be equal intensity at D1 and D2 which will now be proportional to $\rho_{22}$. It can be shown that the average intensity at the two detectors, in the two scenarios, is given by
@@
where we have assumed $\phi=0$. If one is unsure about $\phi$ in the setup, one can tune the phase shifter such that maximum value of $I_{diff}^{2}$ is obtained. In this analysis it is assumed that all intensities are normalized with the total intensity. One may mention here that even if the incoming quanton is not fully coherent, it will not affect results, because each quanton is split into a superposition of the two paths by a unitary action of the beam splitter BS1. Any fluctuation in the phase of the incoming quanton will not affect superposition of the two paths, as that initial phase will factor out of the superposition state. Interference will be affected only by what happens after BS1. Nevertheless, the results may be affected by some experimental factors such as detector efficiency, and other factors depending on the specific setup.
Given that the Mach-Zehnder interferometer with cold neutrons is widely used and there have been significant developments in the field of neutron interferometry [49], the experiment suggested here should be feasible with neutrons, allowing for the measurement of decoherence. The investigation of decoherence through neutron interferometry has been previously suggested [50]. While decoherence in a conventional photonic Mach-Zehnder interferometer in a lab would be negligible, one may use the proposed method to measure decoherence in optical interferometry involving large distances, e.g., as in the setup involving gravitational wave detection. Interferometry with atoms and molecules is a topic of active research, and decoherence is an important issue there [51, 52, 53, 54]. One can think of extending the method proposed here to such matter-wave interferometry experiments.
@@ VI Discussion and conclusions
To summarize, we formulated a measure of decoherence based on the quantification of entanglement of a system with the environment. The measure is given by $\mathcal{D}_{e}=\tfrac{n}{n-1}[1-\text{tr}(\rho^{2})]$ for systems with finite-dimensional Hilbert space, and $\mathcal{D}_{e}=[1-\text{tr}(\rho^{2})]$ for continuous variable systems, where $\rho$ represents the reduced density operator of the system after tracing over the environment degrees of freedom. This decoherence measure has minimum value 0 and maximum value 1. Figuring out the pointer states that emerge out of decoherence of a system is often a difficult issue. In one basis, the off-diagonal elements might diminish more rapidly over time, whereas in another basis, they may also diminish, albeit at a slower rate [44]. The present formulation does not depend on the pointer states. One just needs to evaluate a trace which is basis independent. In our opinion, a strong point of the present measure is that it is easily applicable to a wide variety of situations. On the other hand, QFI doesn’t measure decoherence directly, rather it measures how sensitive a quantum state is to changes in a chosen parameter. Often decrease of QFI may come, not from decoherence, but from some unitary evolution. Without careful modeling, QFI cannot cleanly separate noise from coherent dynamics.
@@
While the entanglement between a system and its environment generally cannot be undone, non-Markovianity can throw up interesting situations, especially in scenarios where the distinction between decoherence and dissipation is blurred [55]. How this decoherence measure holds up in such situations, needs to be explored further.
@@ Acknowledgements.
Author declarations
The authors have no conflicts to disclose.